OpenStudy (anonymous):

How do I find the x intercept and y intercept of each line if x~y = -3

jimthompson5910 (jim_thompson5910):

is it supposed to be x - y = -3 ? or is it supposed to be x + y = -3 ?

OpenStudy (anonymous):

X - y = 3

OpenStudy (anonymous):

Oops x-y =. -3

jimthompson5910 (jim_thompson5910):

To find the x-intercept, you plug in y = 0 and solve for x x - y = -3 x - 0 = -3 x = -3 So the x-intercept is -3, which is really the point (-3, 0) --------------------------------- To find the y-intercept, you plug in x = 0 and solve for y x - y = -3 0 - y = -3 -y = -3 y = -3/(-1) y = 3 So the y-intercept is 3, which is really the point (0,3)